OpenStudy (anonymous):

A system of equations is given below. 2x + 7y = 1 -3x – 4y = 5 Create an equivalent system of equations by replacing the first equation by multiplying the first equation by an integer other than 1, and adding it to the second equation. Use any method to solve the equivalent system of equations (the new first equation with the original second equation). Prove that the solution for the equivalent system is the same as the solution for the original system of equations.

OpenStudy (anonymous):

@midhun.madhu1987

OpenStudy (mokeira):

\[(2x+7y=1)1.5\] to get \[3x+10.5y=1.5\] add the two equations  simplify to get y Take any equation and substitute for y to get x \[2x+7(1)=1\] \[2x=-6\] x=-3

OpenStudy (mokeira):

@Mandre yeah, you are right...thanks!

OpenStudy (mandre):

@Mokeira That won't do as the question states you have to multiply the 1st equation by an INTEGER other than 1 and add it to the 2nd equation. I removed my calculations as I made a mistake. You are supposed to add the 2nd to the first and I went vice versa.